OpenStudy (anonymous):

The ratio of the measures of two supplementary angles is 7:8. Which proportion(s) could you use to find the measures of the angles?

OpenStudy (anonymous):

This is multiple choice, some of the answers are: x/y= 7/8 x/y=8/7 x/180-x=7/8 x/180-x=8/7

OpenStudy (anonymous):

@HWBUSTER00

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

try both 3rd or 4th

OpenStudy (anonymous):

Thanks for replying, how did you come to this conclusion?

OpenStudy (anonymous):

\[x+y=180,y=180-x\] \[\frac{ x }{ y }=\frac{ x }{ 180-x }=\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Thank you very much for your help!